Suppose you are a U.S. investor who is planning to invest $315,000 in Japan. You do so at a starting exchange rate of 86.18 ¥/$. Your Japanese investment gains 8.90 percent, and the ending exchange rate is 84.46 ¥/$. What is your total return on this investment?
Answer #1
Initial Investment in USD = 315,000
Initial Exchange rate = 86.18 ¥/$.
1$ = 86.18 ¥
$315000 = ¥ 86.18*315,000 = ¥ 27,146,700
Initial Investment in Yen = ¥ 27,146,700
Gain = 8.9%
Gain = 8.9% * 27,146,700 = ¥ 2,416,056.30
Final investment in Yen = Initial Investment in Yen + Gain
Final investment in Yen = ¥ 27,146,700 + ¥ 2,416,056.30
Final investment in Yen = ¥ 27,146,700 + ¥ 2,416,056.30 = ¥ 29,562,756.30
Final exchange rate = 84.46 ¥/$
I$ = 84.46 ¥
1 ¥ = 1/84.46 $
Final investment in USD = $ 29,562,756.30 * 1/84.46 = $ 350,020.7945
Return on investment = (Final investment/Initial Investment -1 ) * 100
Return on investment = (350,020.7945/315,000 -1 ) * 100
Return on investment = 11.11771253%
Return on investment = 11.12% (approx)